OpenStudy (anonymous):

The length of pregnancies are normally distributed with a mean of 268 days and a standard deviation of 15 days. Find the probability of a pregnancy lasting 280 days or longer

OpenStudy (kirbykirby):

Let \(X\) be the number of days a pregnancy lasts. The question asks you for P\((X\ge 280)\). WHat you must do is standardize \(X\) and then you'll be able to use a standard normal table: \[P(X\ge 280)=P\left( \frac{X-268}{15}\ge\frac{280-268}{15}\right)=P(Z\ge0.8) \]